Evening Shift 23/18/99

Peter Shanahan (shanahan@fnal.gov)
Wed, 24 Nov 1999 00:18:07 -0600

Shift Summary for eve shift, 23rd Nov 99

On Shift: PSH, Wah

Hours of beam available at all intensities: 0.16
Hours of beam available at nominal intensity: 0.00
Hours of E799 data (special runs not included): 0.00
Hours of special runs (laser scans, muon runs...): 1.00

Accounting of lost time:
Shutdown period

========================================================================
Runs ended this shift:

Run 15037 - - LASER SCAN
Run 15039 - 0.3M B01 - BEAM_MUON

Tapes brought to FCC this shift:
(including tapes for online split)

none

========================================================================
Accesses during shift:

Several to investigate VV' latch problem

========================================================================
Overview of shift:
Recover from 1 day shutdown

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15:50 JW re-establishes real spill signal.
16:30 JRL changes cabling on 3X KQ to fix trigger
hanging. Did it work....?
17:05 2 new Ar-Ethane bottles on line (DN).
17:10 JW installed new CSQ6 constants
18:00 VJ and SL tweak crate voltages - no effect on VV'
19:05 RJT loosened MU3 Alarm limits
19:11 Hamm steals UPX013 for super secret analysis.
19:42 PSH changes laser delay to 10ns from 11ns to recover
from slow drift since run 15000
20:00 BobT, SL, VJ spend more time investigating VV' latches
21:15 "Startup in progress...."
15037 1 22:21 Start laser scan
22:35 Raise vacuum shutter
15039 1 23:47 Beam returns, start muon run - 2E12 - 5E12
23:57 Request nominal magnet configuration

========================================================================
Follow up Notes on Previous Shift

- init problems in pipeline were due to triggers during init.

========================================================================
Items to Follow up on Next Shift

- download new q/e constants once beam has been back for ~15'

========================================================================
Items for Experts and Management